OpenStudy (anonymous):

4-z 2 --- = --- solve z= 6-z z-6

OpenStudy (anonymous):

(4-z)(z-4) = (2)(6-z) 4z - 16 -z^2 +4z = 12 - 2z -z^2 +8z - 16 = 12 - 2z -z^2 +10z -28 = 0 z^2 - 10z + 28 = 0 (z-6 ) (z-4) = 0 equating the 2 factors by 0, we have : z - 6 =0 ; z -4 =0 z = 6 ; z = 4. :]]